OpenStudy (anonymous):

In a football game a kicker attempts a field goal. The ball remains in contact with the kicker's foot for 0.0800 s, during which time it experiences an acceleration of 229 m/s2. The ball is launched at an angle of 52.3 ° above the ground. Determine the (a) horizontal and (b) vertical components of the launch velocity.

OpenStudy (anonymous):

launch velocity = a * t = 229 *0 .08 m/s vertical component = (229*0.08) * sin 52.3 horizontal component = (229*0.08) * cos 52.3

OpenStudy (anonymous):

need any more help?

OpenStudy (anonymous):

I have some other questions. Feel like a moron, though. It's been a while since I've looked at stuff like this. :P

OpenStudy (anonymous):

unfortunately i need to go in 5 minutes do you want me to work out the above computations?

OpenStudy (anonymous):

If you could.

OpenStudy (anonymous):

Much appreciated.

OpenStudy (anonymous):

launch velocity - 18.32 m/s vertical component = 14.5 m/s horizontal component = 11.2 m/s